OpenStudy (anonymous):

What is the limit for x--->infinity (r+sinr) / (2r+7-5sinr)

OpenStudy (anonymous):

I guess you mean as r tend to infinity? Divide the top and bottom of the fraction both by r. Then use the fact: \[\lim_{r \rightarrow \infty} \frac{\sin r}{r} = 0\](since sin(x) remains bounded between -1 and 1, whereas r goes off to infinity.)

OpenStudy (anonymous):

Yep, woops. Thank you

OpenStudy (anonymous):

You're welcome :)

OpenStudy (rob1525):

Um both the top and bottom go to infinity...so...the limit shouldn't exist. let me verify with my calc.

OpenStudy (rob1525):

Nope im getting 1/2. Humm.....

OpenStudy (anonymous):

@robtobey Remember in maths \(\frac{\infty}{\infty}\) could be anything - it's basically meaningless. Just because the top of the fraction tends to infinity, and the bottom tends to infinity, does NOT mean the limit doesn't exist!

OpenStudy (anonymous):

(Sorry, quoted wrong person!) @rob1525 See above.

OpenStudy (rob1525):

oh, geees. if the sin are eliminated then ur left with x/2x which is a horizontal retricemtote of 1/2.

OpenStudy (rob1525):

*asymptote